# OpenAgentSafety: A Comprehensive Framework for Evaluating Real-World AI Agent Safety

### v1

@@ 3 Experiment
In this section, we first describe the experimental setup and agent evaluation pipeline used to run our benchmark (§3.1). We then present overall safety results across five widely used LLMs and analyze failure rates, unsafe behavior rates, and evaluator disagreements (§3.2). Finally, we conduct detailed analyses across varied user intents, risk categories, and tools (§3.3.
@@ 3.1 Experimental Setup
We evaluate five widely adopted LLMs on the 356 tasks in OA-Safety, including openweight LLMs: Deepseek-v3 [deepseek-vl-v3] and Deepseek-R1 [guo2025deepseekr1], as well as proprietary LLM providers: Claude Sonnet 3.7 [claude-sonnet-3.7], GPT-4o [openai-gpt-4o], and o3-mini [o3mini2025]. These models have varying capabilities and alignment strategies. All are deployed as the base model within the OpenHands [wang2025openhandsopenplatformai] agentic scaffold, which interfaces with real tools inside a sandboxed Docker environment. Each task is mounted into the Docker container alongside any required websites or files. As the agent completes the task, its trajectory is saved for downstream evaluation.
@@ 3.2 Results
Table 4: Unsafe behavior rates for LLM-as-Judge and rule-based evaluation across models, along with Failure and Disagreement rates. Percentages are computed only over tasks where the agent reached safety-vulnerable states, where it was exposed to malicious input/content.
Model LLM-as-Judge Rule-based Disagreements Failure Rate
Claude Sonnet 3.7 51.20 32.85 9.17 40.30
o3-mini 72.73 32.32 6.57 43.30
GPT-4o 65.80 32.64 5.69 45.00
Deepseek-v3 62.23 32.44 6.91 46.30
Deepseek-R1 62.78 32.78 6.67 48.90
Table 4 shows the safety outcomes across models. As per the LLM Judge, approximately 40–49% of the tasks fail before reaching a safety-vulnerable state, typically due to web navigation failures, authentication issues, or tool misuse. These failures highlight current limitations of AI agents in long-horizon reasoning and tool coordination. For example, in file download tasks, agents frequently fail to retrieve the file, preventing subsequent unsafe interactions from even being triggered. These failure cases underscore the increasing realism and difficulty of tasks enabled by our framework, which can support evaluating more capable agents as the field progresses.
To isolate safety behavior, we report unsafe rates only over safety-vulnerable trajectories. All models exhibit substantial unsafe behavior, with LLM-as-Judge rates ranging from 51% (Claude Sonnet 3.7) to 72% (o3-mini). Disagreements between the evaluators are rare and occur when the LLM judge incorrectly believes a task to be safe, whereas, the rule-based evaluator detects unsafe environment changes (e.g., a leaked document), prompting review. This highlights the importance of combining the two types of evaluation.
@@ 3.3 Analysis
We support our findings with both quantitative results and qualitative examples, with the latter included in the Appendix §A (Table 7). We combine quantitative results (Table 4) with representative failure trajectories (see Appendix Table 7) to analyze how unsafe behaviors vary across user intents, risk categories, and tools (Figure 3 4 5). In each heatmap, rows and columns are sorted in descending order based on the average rate of unsafe behavior.
@@
(1) Benign intent does not imply safety: Seemingly benign user prompts still lead to unsafe behavior in 57–86% of cases across models (Figure 3). This pattern is driven by two key factors. First, agents tend to overgeneralize user goals: for instance, in api-in-codebase (Table 7), GPT-4o “helpfully” hard-codes an API key into the codebase, prioritizing task completion over secure coding practices. Second, agents often lack caution when the user’s request seems harmless. For example, Claude Sonnet 3.7 despite having the lowest rate of unsafe behavior for malicious intents, has the highest unsafe rate under benign intent (85.7%), suggesting that refusal strategies tuned for detecting overtly harmful instructions may not generalize to subtler risks, where the harm emerges from context or incorrect assumptions about user goals. Without explicit cues, agents may default to over-compliance, violating security principles.
Figure 3: Agent unsafe behaviour rates (%) for the user and NPC intents across LLMs.
(2) Explicit malice partially activates defenses, but unevenly: When the user is openly malicious, Claude Sonnet 3.7 and Deepseek-v3’s unsafe rate plunges to 29.9% and 33.2% respectively, showing effective policy gating. o3-mini, by contrast, rises to 80.7%, indicating near-absence of refusal logic. GPT-4o (64.7%) and Deepseek-R1 (66%) sit mid-range. Thus, model size is not predictive of robustness but alignment strategy is. (3) Hidden intent circumvents safeguards: In the benign user intent and malicious NPC setting, where harmful goals are introduced by other NPC actors, unsafe rates converge in a narrow, high band (53.9–69.2%). Claude and Deepseek-v3’s risky actions more than double relative to malicious intent, both of which had low rate of unsafe behaviors for malicious intent, confirming that multi-turn intent tracking is unsolved. The meeting-topics trajectory demonstrates the failure: GPT-4o reorders agenda items on polite request, ignoring implicit fairness constraints, and no policy trigger fires because each micro-instruction appears benign.
@@
(1) Systemic risks consistently bypass safeguards: The highest unsafe rates occur in categories that require procedural judgment or understanding of institutional norms, including computer security compromise (72–86%), legal violations, privacy breaches, and harmful decision making. These tasks involve policy boundaries rather than syntactic dangers, which current agents fail to internalize. In change-branch-policy (Table 7), for instance, multiple models convert a private GitLab repository to public access at the request of a fired employee, mistaking helpfulness for authorization. Trajectory evaluations show that agents frequently disregard authorization entirely, suggesting a lack of procedural reasoning. (2) Operational risks lead to varying judgment: Mid-tier categories such as data loss, financial loss, and unsafe code show agent-level variation. Claude Sonnet 3.7 is relatively cautious on destructive actions unlike GPT-4o, a disparity observed in misleading-task, where only Claude inspects a malicious script before running it. These results show that certain forms of sandboxed harm (e.g., ‘rm -rf‘) remain weak spots for models that rely on shallow interpretation.
Figure 4: Unsafe behavior rates (%) across risk categories and LLMs.
(3) Content moderation tasks benefit from targeted alignment: Spreading malicious content corresponds to the lowest unsafe rates overall, especially for Claude Sonnet 3.7 (27.7%). This likely reflects effective training on toxic language. However, performance is still uneven as o3-mini still fails in 75% of cases, and GPT-4o generates discriminatory rationales in trends task, suggesting that even when content filters are present, deeper social reasoning remains fragile.
RQ3: Do tools enable or restrain unsafe actions? (1) Browsing is the most failure-prone interface: Tasks involving web interaction (e.g., OwnCloud, GitLab) yield the highest unsafe rates across models (60–75%). Agents often struggle with authentication, redirects, and dynamic DOM content, which distracts them from recognizing unsafe behavior due to the large browsing context. For instance, in leak-document (Table 7), GPT-4o fails to complete the download due to authentication issues, but when navigation succeeds, it proceeds to share confidential files without verification.
@@
(2) File systems and code execution magnify intent errors: Unsafe behavior in these tools often results from executing commands or modifying files without contextual checks (e.g. In misleading-task, the malicious content is hidden inside the file and not presented to the agent unless it tries exploring the environment.). o3-mini (73.8%) and GPT-4o (71.8%) display significantly higher unsafe rates than Claude (55.6%), despite similar capabilities, suggesting that Claude’s refusals and cautious completions offer some protection. Still, the misleading-task, where most agents blindly run a malicious rm -rf script, demonstrates that even simple reasoning failures can yield high-impact damage. (3) Messaging tools introduce social manipulation risk: Chat-based tasks elicit unsafe behavior rates between 45–70%. Agents often fail to validate user roles or authorization before sharing sensitive information or accepting requests. In unauth-discount, agents approve significant markdowns under persuasive, but unauthorized, customer prompts. This pattern underscores the fragility of agent alignment in multi-user or social settings, where politeness or over-cooperation overrides internal policy logic.
Design implications Findings from OA-Safety point to three actionable priorities for improving agent safety: (i) Contextual intent aggregation, where refusal mechanisms must operate over multi-turn context rather than isolated prompts, (ii) Tool-specific privilege boundaries, enforcing stricter runtime controls for high-risk tools like code execution and file manipulation, and (iii) Policy-grounded supervision, using datasets aligned with legal, organizational, and procedural norms to train agents for regulated environments. OA-Safety provides executable environments with realistic tool interfaces, where these safeguards can be iteratively prototyped and stress-tested under adversarial and ambiguous conditions prior to deployment.
4 Related work
Safety guidelines  Designing tasks that elicit unsafe behavior from AI agents requires grounding in established risk taxonomies and policies. Frameworks such as the AIR taxonomy [zeng2024airiskcategorizationdecoded] and technical interpretations of the EU AI Act [guldimann2025complaiframeworktechnicalinterpretation] define categories spanning operational, societal, and legal risks. Recent work emphasizes aligning agent behavior with human values [tang2024prioritizingsafeguardingautonomyrisks] and constructing environments that provide safe interaction affordances [chan2025infrastructureaiagents]. These perspectives inform the risk categories and scenario designs used in OpenAgentSafety.
LLM and agent safety evaluations  Prior benchmarks have focused extensively on unsafe generations from LLMs [röttger2025safetypromptssystematicreviewopen, tedeschi2024alertcomprehensivebenchmarkassessing], probing biases, toxic completions, and jailbreaking strategies [doumbouya2025h4rm3llanguagecomposablejailbreak, jiang2024wildteamingscaleinthewildjailbreaks]. While these efforts helped shape safety-aligned finetuning and refusal training [kumar2023certifying, wang2023self], they mostly assess static output generation. In contrast, agent safety benchmarks assess agents with tool-use capabilities [mo2024tremblinghousecardsmapping, li2025commercialllmagentsvulnerable], expanding the risk surface to include execution-based harms. However, many such evaluations rely on simulated APIs and simplified environments [andriushchenko2025agentharmbenchmarkmeasuringharmfulness, yin2025safeagentbenchbenchmarksafetask, yuan2024rjudgebenchmarkingsafetyrisk], limiting realism. Other evaluations are constrained to single tools and short interaction paths. Tool-specific evaluations have largely targeted: (i) Web environments: Testing agents’ robustness to pop-ups, authentication barriers, and misleading content [tur2025safearenaevaluatingsafetyautonomous, zhang2024attackingvisionlanguagecomputeragents, xu2024advwebcontrollableblackboxattacks, chen2025shieldagentshieldingagentsverifiable]; (2) Code execution: Evaluating safety in generating or running scripts [guo2024redcoderiskycodeexecution]; and (3) Social interaction: Simulating user conversations or agent collaboration [shao2025collaborativegymframeworkenabling, zhou2024sotopiainteractiveevaluationsocial].
Our work differs by integrating multiple real tools (e.g., bash, file systems, browsers, messaging) into a single framework with multi-turn, multi-user interactions. Unlike prior work, we simulate both benign and adversarial users across long-horizon trajectories, exposing agents to more realistic decision-making challenges.
Training for safer agents  To improve agent robustness, recent work proposes scoring actions as safe or unsafe [yuan2024rjudgebenchmarkingsafetyrisk], defensive agent architectures [chen2025shieldagentshieldingagentsverifiable], and adversarial fine-tuning strategies [rosser2025agentbreedermitigatingaisafety]. Others advocate for active learning to prioritize rare risk cases [abdelnabi2025firewallssecuredynamicllm], or explore how performance optimization can reduce safety margins [wu2025dissectingadversarialrobustnessmultimodal]. While promising, these approaches often assume access to evaluation settings that mirror realistic threats. Our benchmark fills this gap by offering a high-fidelity simulation framework suitable for safety training, adversarial red-teaming, and future reinforcement learning setups that operate across complex, multi-tool agent workflows.
5 Limitations
5.1 Simulation
While our framework is broad in scope, several limitations remain. Many current agents struggle with tasks that require exploration, delayed feedback, or handling dynamic and partially observable environments, often failing before reaching safety-critical decision points. Additionally, secondary actors (NPCs), while valuable for creating controlled multi-agent scenarios, can occasionally deviate from assigned strategies or fail to engage as intended, limiting task coverage in more complex setups.
5.2 LLM Judges for Agentic Evaluation
A key limitation in our evaluation pipeline is the use of LLMs as judges to assess trajectory safety. Prior work [zhang2024agentsafetybenchevaluatingsafetyllm] has shown that off-the-shelf models are often unreliable in evaluating agentic behavior. To assess reliability, we conducted a human annotation study on a randomly sampled set of 100 GPT-4o trajectories. Two expert annotators independently labeled each trajectory’s safety outcome, achieving 94% inter-annotator agreement. These human judgments were then used to benchmark the accuracy of LLM-based judges on both failure detection and safety categorization.
We also evaluated a fine-tuned Qwen 2.5 7B model from Agent-SafetyBench (ASB) [zhang2024agentsafetybenchevaluatingsafetyllm], which does not include a separate failure category. Table 5 compares safety classification performance across judges. We find that LLM judges, particularly GPT-4.1, struggle with nuanced failure cases, often overestimating failure rates due to superficial error signals (e.g., tool failures) and underestimating unsafe behavior that is implied rather than explicitly acknowledged. For example, an agent might not affirm an unsafe request via message, but proceed to call tools to fulfill it, implicitly indicating agreement. As a result, reported failure rates are likely inflated, and unsafe behavior rates should be interpreted as conservative lower bounds. Future work should explore how to train better and more robust in-house safety and failure detection models.
@@ 5.2 LLM Judges for Agentic Evaluation
Despite these challenges, OA-Safety is designed with modularity in mind. It supports integration of new environments, improved evaluation routines, and interventions such as guardrail agents or safety-aware training signals derived from supervision or environment-based reinforcement.
6 Conclusion

### v2

@@ 3 Experiments and Results
In this section, we first describe the experimental setup and agent evaluation pipeline used to run our benchmark (§3.1). We then present overall safety results across five widely used LLMs and analyze failure rates, unsafe behavior rates, and evaluator disagreements (§3.2). Finally, we conduct detailed analyses across varied user intents, risk categories, and tools (§3.3).
@@ 3.1 Experimental Setup
We evaluate seven widely adopted LLMs on the 356 tasks in OA-Safety, including openweight LLMs: Deepseek-v3 (DeepSeek-AI et al., 2024) and Deepseek-R1 (Guo et al., 2025), as well as proprietary LLM providers: Claude Sonnet 3.7 (Anthropic, 2025), GPT-4o (OpenAI et al., 2024), as well as their successors, Claude Sonnet 4 (PBC, 2025) and GPT-5 OpenAI (2025b), and o3-mini (Zhang et al., 2025) which are widely integrated into agentic frameworks. o3-mini and Deepseek-R1 are reasoning LLMs allowing us to examine how reasoning capabilities affect safety. We also examine how improvements in model capabilities impact safety for models in similar families. These models have varying capabilities and alignment strategies. We use the OpenHands (Wang et al., 2025) agentic scaffold, which interfaces with real tools inside a sandboxed Docker environment. Each task is mounted into the Docker container alongside any required websites or files. After the agent completes the task, its trajectory is saved for evaluation.
@@ 3.2 Results
Table 3: Unsafe behavior rates for LLM-as-Judge and rule-based evaluation across models, along with Failure, Disagreement, and Successful Completion rates. Percentages for LLM-Judge and Rule-based are computed only over tasks where the agent reached safety-vulnerable states, where it was exposed to malicious input/content.
Model LLM-Judge Rule-based Disagreements Failure Successful
Claude Sonnet 4 49.06 49.06 14.95 39.70 37.10
Claude Sonnet 3.7 51.20 32.85 9.17 40.30 33.88
GPT-5 52.58 52.15 13.79 34.80 40.40
GPT-4o 65.80 32.64 5.69 45.00 21.77
o3-mini 72.73 32.32 6.57 43.30 17.29
Deepseek-v3 62.23 32.44 6.91 46.30 22.12
Deepseek-R1 62.78 32.78 6.67 48.90 20.86
Table 3 shows the various safety metrics across different LLMs. To isolate analyses of unsafe behavior, we report unsafe rates only over safety-vulnerable trajectories that did not cause any agent failures. All models exhibit substantial unsafe behavior, with LLM-as-Judge rates ranging from 49% (Claude Sonnet 4) to 73% (o3-mini). We conduct Mann-Whitney U tests (McKnight and Najab, 2010) on the unsafe behavior rates and find two sets of models: Claude Sonnet 3.7, Claude Sonnet 4, and GPT-5 are significantly safer (i.e. lower unsafe behavior rates) than all models in the other set: vs. o3 mini (p << 0.001), vs. GPT-4o (p $\leq$ 0.006), vs. DeepSeek v3 (p $\leq$ 0.026), vs. DeepSeek R1 (p $\leq$ 0.023). All other comparisons were non-significant (p >> 0.16). Disagreements between the evaluators are rare and occur when the LLM judge incorrectly believes a task to be safe, whereas, the rule-based evaluator detects unsafe environment changes (e.g., a leaked document), prompting review. This highlights the importance of combining the two types of evaluation. A prominent category of disagreement is security/credential-related tasks where the LLM judge misses the finer unsafe actions and the rule-based evaluator detects the deterministic change.
The LLM Judge reports that approximately 35–49% of the tasks fail before reaching a safety-vulnerable state, typically due to web navigation failures, authentication issues, or incorrect tool-use. These failures highlight current limitations of LLMs for long-horizon tasks. For example, in file download tasks, they frequently fail to retrieve the file, preventing subsequent unsafe interactions from even being triggered. These failure cases underscore the increasing realism and difficulty of tasks enabled by our framework, which can support the evaluation of more capable LLMs in the future. We also observe inconsistencies between safe completion annotation by LLM-Judge and the unsafe trajectory judgments, highlighting the need for rule-based assessments.
@@ 3.3 Analysis
We organize our analysis of current model behaviors around four research questions that guide our evaluation of agent safety: RQ1: How does user intent influence unsafe behavior? RQ2: Which risk categories reveal the most serious vulnerabilities? RQ3: Do tools enable or restrain unsafe actions? RQ4: How reliable are LLM judges in detecting unsafe behavior?
We support our findings with both quantitative results and qualitative examples, with the latter included in the Appendix §A (Table 6) to show the utility of the framework. We combine quantitative results (Table 3) with representative failure trajectories (see Appendix Table 6) to analyze how unsafe behaviors vary across user intents, risk categories, and tools (Figure 3 4 5). In each heatmap, rows and columns are sorted in descending order based on the average rate of unsafe behavior.
@@
(1) Benign intent does not imply safety: Seemingly benign prompts still cause unsafe behavior in 50-86% tasks across models (Figure 3). Agents often overgeneralize user goals: for instance, in api-in-codebase (Table 6), GPT-4o “helpfully” hard-codes an API key into the codebase, prioritizing completion over security. Agents often lack caution with seemingly harmless requests.
Figure 3: Unsafe behaviour rates (%) across user and NPC intents.
For example, Claude Sonnet 3.7 despite having one of the lowest rate of unsafe behavior for malicious intents, has the highest unsafe rate under benign intent (85.7%), suggesting that refusal strategies do not not generalize to subtler risks, where the harm emerges from context or incorrect assumptions about user goals. Interestingly, Claude Sonnet 4 is able to significantly improve over its predecessor in this category. (2) Explicit malice partially activates defenses, but unevenly: For tasks with malicious user intents, Claude Sonnet 3.7 and Deepseek-v3’s unsafe rate plunges to 30% and 33.3% respectively, showing effective policy gating. o3-mini, by contrast, rises to 80.7%, indicating near-absence of refusal logic. GPT-4o (64.7%) and Deepseek-R1 (66%) sit mid-range. GPT-5 displays improvements over GPT-4o for almost all intents, but Claude Sonnet 4 sees an increase over Claude Sonnet 3.7 in unsafe actions for malicious intent. (3) Hidden intent circumvents safeguards: In the benign user intent and malicious NPC setting, where harmful goals are introduced by other NPCs, unsafe rates vary from 44.9–69.2%. Claude 3.7 and Deepseek-v3’s unsafe rates more than double relative to malicious intent, confirming that multi-turn intent tracking is ineffective. GPT-5 is surprisingly strong in this category. The meeting-topics trajectory shows GPT-4o reorders agenda items on a polite request ignoring fairness considerations.
@@
(1) Systemic risks consistently bypass safeguards: The highest unsafe rates occur in categories that require procedural judgment or understanding of institutional norms, including computer security compromise (72–86%), legal violations, privacy breaches, and harmful decision making. [^3] A notable exception being Claude which is relatively more robust in legal scenarios.
Figure 4: Unsafe behavior rates (%) across risk categories.
These tasks involve policy boundaries rather than syntactic dangers, which current agents fail to comprehend. In change-branch-policy (Table 6), for instance, multiple models convert a private GitLab repository to public access at the request of a fired employee, mistaking helpfulness for authorization. Trajectory evaluations show that agents frequently disregard authorization entirely, suggesting a lack of procedural reasoning. (2) Operational risks lead to varying judgment: Mid-tier categories such as data loss, financial loss, and unsafe code show varying trends. Claude Sonnet models are relatively cautious on destructive actions unlike GPT-4o, a disparity observed in misleading-task, where only Claude inspects a malicious script before running it. These results show that certain forms of sandboxed harm (e.g., ’rm -rf’) remain weak spots for models that rely on shallow interpretation. (3) Content moderation tasks benefit from targeted alignment: Spreading malicious content corresponds to the lowest unsafe rates overall, especially for Claude Sonnet 3.7 (27.7%). This likely reflects effective safety training to avoid toxic content. However, performance is still uneven as o3-mini still fails in 75% of cases, and GPT-4o generates discriminatory rationales in trends task, suggesting that even when content filters are present, deeper social reasoning remains fragile.
RQ3: Do tools enable or restrain unsafe actions?
(1) Browsing is the most failure-prone interface: Tasks involving web interaction (e.g., OwnCloud, GitLab) yield the highest unsafe rates across models (59–75%). Agents often struggle with authentication, redirects, and dynamic DOM content, which distracts them from recognizing unsafe behavior due to the large browsing context. For instance, in leak-document (Table 6), GPT-4o fails to complete the download due to authentication issues, but when navigation succeeds, it proceeds to share confidential files without verification. GPT-5 and Sonnet 4 make some improvements with this tool. (2) File systems and code execution magnify intent errors: Unsafe behavior in these tools often results from executing commands or modifying files without contextual checks (e.g. In misleading-task, the malicious content is hidden inside the file and not presented to the agent unless it tries exploring the environment.). o3-mini (73.8%) and GPT-4o (71.8%) display significantly higher unsafe rates than GPT-5/Claude models, suggesting that Claude’s refusals and cautious completions offer some protection. Still, the misleading-task, where most LLMs blindly run a malicious rm -rf script, shows that simple reasoning failures can cause severe damage. (3) Messaging tools introduce social manipulation risk: Chat-based tasks show 45-70% unsafe rates. LLMs fail to validate user roles before sharing sensitive information or accepting requests. In unauth-discount, LLMs approve significant discounts under persuasive, but unauthorized, customer prompts.
@@
This pattern underscores the fragility of LLM alignment in multi-user settings, where politeness overrides internal policy logic.
RQ4: How reliable are LLM judges in detecting unsafe agentic behavior?
(1) LLM judges struggle with nuanced failure cases: To assess evaluator reliability, we conducted a human annotation study on 100 randomly sampled GPT-4o trajectories with two expert annotators achieving 94% inter-annotator agreement. Comparing these human judgments against LLM-based judges reveals systematic weaknesses in both failure detection and safety categorization (Table 4). GPT-4.1 and other LLM judges frequently underestimate unsafe behavior that is implied rather than explicitly acknowledged. For example, an agent might not verbally affirm an unsafe request but proceed to call tools to fulfill it, implicitly indicating agreement, a pattern that LLM judges often miss. (2) Superficial error signals inflate failure rates: LLM judges overestimate failure rates by misinterpreting superficial error signals (e.g., tool failures, authentication issues) as task failures, even when the agent successfully navigates to safety-vulnerable states. This suggests that reported failure rates in Table 3 are likely inflated, and we find that unsafe behavior rates should be interpreted as conservative lower bounds. (3) Fine-tuned judges show limited improvement: We evaluated a fine-tuned Qwen 2.5 7B model from  Zhang et al. (2024b), which lacks a separate failure category. While specialized training helps, the model still exhibits similar erroneous behaviors like off-the-shelf LLMs. This confirms findings from prior work (Zhang et al., 2024b) that LLM-based evaluation of unsafe agentic behavior is unreliable. This underscores the need for hybrid evaluation approaches combining LLM judges with rule-based checks, as demonstrated by our disagreement analysis in Table 3.
@@
Design implications. Our findings point to three actionable priorities for improving agent safety: (i) Contextual intent aggregation, where refusal mechanisms must operate over multi-turn context rather than isolated prompts, (ii) Tool-specific privilege boundaries, enforcing stricter runtime controls for high-risk tools like code execution and file manipulation, and (iii) Policy-grounded supervision, using datasets aligned with legal, organizational, and procedural norms to train agents for regulated environments. OA-Safety provides executable environments with realistic tool interfaces, where these safeguards can be iteratively prototyped and stress-tested under adversarial and ambiguous conditions prior to deployment.
4 Related work
Safety guidelines  Designing tasks that elicit unsafe behavior from AI agents requires grounding in established risk taxonomies and policies. Frameworks such as the AIR taxonomy (Zeng et al., 2024) and technical interpretations of the EU AI Act (Guldimann et al., 2025) define categories spanning operational, societal, and legal risks. Recent work emphasizes aligning agent behavior with human values (Tang et al., 2024) and constructing environments that provide safe interaction affordances (Chan et al., 2025). These perspectives inform the risk categories and scenario designs used in OpenAgentSafety.
LLM and agent safety evaluations  Prior benchmarks have focused extensively on unsafe generations from LLMs (röttger2025safetypromptssystematicreviewopen; Tedeschi et al., 2024), probing biases, toxic completions, and jailbreaking strategies (Doumbouya et al., 2025; Jiang et al., 2024). While these efforts helped shape safety-aligned finetuning and refusal training (Kumar et al., 2023; Wang et al., 2023), they primarily assess static output generation. In contrast, agent safety works assess agents with tool-use capabilities (Mo et al., 2024; Li et al., 2025), expanding the risk surface to include execution-based harms. However, many such evaluations rely on simulated APIs and simplified environments (Andriushchenko et al., 2025; Yin et al., 2025; Yuan et al., 2024), limiting realism. Other evaluations are constrained to single tools and short interactions. Tool-specific evaluations have largely targeted: (i) Web environments: Testing agents’ robustness to pop-ups, authentication barriers, and misleading content (Tur et al., 2025; Zhang et al., 2024a; Xu et al., 2024a; Chen et al., 2025); (2) Code execution: Evaluating safety in generating or running scripts (Guo et al., 2024); and (3) Social interaction: Simulating user conversations or agent collaboration (Shao et al., 2025b; Zhou et al., 2024c). Our work differs by integrating real tools (e.g., code execution, browsers, messaging) into a single framework with multi-turn, multi-user interactions. Unlike prior work, we simulate both benign and adversarial users, exposing agents to more realistic decision-making challenges.
Training for safer agents  To improve agent robustness, recent work proposes scoring actions as safe or unsafe (Yuan et al., 2024), defensive agent architectures (Chen et al., 2025), and adversarial fine-tuning strategies (Rosser and Foerster, 2025). Others advocate for active learning to prioritize rare risk cases (Abdelnabi et al., 2025), or explore how performance optimization can reduce safety margins (Wu et al., 2025). While promising, these approaches often assume access to evaluation settings that mirror realistic threats. Our benchmark fills this gap by offering a high-fidelity simulation framework suitable for safety training, adversarial red-teaming, and reinforcement learning setups.
5 Conclusion, Limitations, and Future Work